# Neural-Driven Image Editing

### v2

@@ Abstract
Traditional image editing typically relies on manual prompting, making it labor-intensive and inaccessible to individuals with limited motor control or language abilities. Leveraging recent advances in brain-computer interfaces (BCIs) and generative models, we propose LoongX, a hands-free image editing approach driven by multimodal neurophysiological signals. LoongX utilizes state-of-the-art diffusion models trained on a comprehensive dataset of 23,928 image editing pairs, each paired with synchronized electroencephalography (EEG), functional near-infrared spectroscopy (fNIRS), photoplethysmography (PPG), and head motion signals that capture user intent. To effectively address the heterogeneity of these signals, LoongX integrates two key modules. The cross-scale state space (CS3) module encodes informative modality-specific features. The dynamic gated fusion (DGF) module further aggregates these features into a unified latent space, which is then aligned with edit semantics via fine-tuning on a diffusion transformer (DiT). Additionally, we pre-train the encoders using contrastive learning to align cognitive states with semantic intentions from embedded natural language. Extensive experiments demonstrate that LoongX achieves performance comparable to text-driven methods (CLIP-I: 0.6605 vs. 0.6558; DINO: 0.4812 vs. 0.4636) and outperforms them when neural signals are combined with speech (CLIP-T: 0.2588 vs. 0.2549). These results highlight the promise of neural-driven generative models in enabling accessible, intuitive image editing and open new directions for cognitive-driven creative technologies. Datasets and code will be released to support future work and foster progress in this emerging area.
@@ 1 Introduction
Image editing involves manipulating digital visuals to achieve desired effects, significantly impacting fields like advertising, entertainment, and scientific visualization [1]. Traditionally, this task demands extensive manual effort and technical expertise. Advances in generative models have streamlined instruction-based image editing through automated pipelines [2, 3, 4]. Nevertheless, these methods still heavily depend on intensive user inputs, such as text prompts [5, 6], visual references like masks or sketches [7, 8], and physical operations like dragging [9, 10, 11, 12]. Such reliance limits efficiency and accessibility, especially for users with motor or communication impairments.
To address these challenges, alternative input modalities have been explored [13, 14, 15] for image editing. Among these, brain–computer interfaces (BCIs) provide a promising possibility with their recent advancement in hardware precision [16, 17]. Starting from early attempts in passive tasks such as mental state recognition [18] and neural activity analysis [19, 20], BCIs have begun to be involved in more active generative tasks such as neural-driven chat [21] and visual content creation [22, 23].
However, existing approaches remain limited to single-modality data such as electroencephalography (EEG) [24, 22] or functional magnetic resonance imaging (fMRI) [25], which is insufficient to capture nuanced user intentions for complex editing scenarios. In practice, physiological signals from different modalities offer complementary insights into cognitive states such as attention, motivation, and emotional regulation [26, 27, 28, 29, 30, 31, 32], underscoring the need for multimodal neural integration.
@@
To answer these questions, we construct L-Mind, a comprehensive multimodal dataset comprising 23,928 image pairs with synchronously collected EEG, functional near infrared spectroscopy (fNIRS) [33], photoplethysmography (PPG) [34], head motion, and speech signals from 12 participants conceiving image editing tasks. Captured using a wireless, lightweight BCI system that supports unconstrained head movements and speech [35], L-Mind offers higher ecological validity under natural real-world conditions and supports robust training of brain-supervised generative models.
Building on L-Mind, we also propose LoongX, a hands-free image editing approach that innovatively integrates the proposed multimodal neural signal fusion strategy with a diffusion transformer (DiT) to translate neural intent into image edits. Unlike prior single-modal methods, LoongX integrates EEG, fNIRS, PPG, and head motion signals, extracting explicit user intentions from EEG signals across multiple scalp regions, incorporating cognitive load and emotional valence data from fNIRS, and capturing stress and engagement indicators through PPG and motion signals. We introduce two new modules to manage diverse multimodal input: a cross-scale state space (CS3) encoder for robust feature extraction and a dynamic gated fusion (DGF) module for comprehensive multimodal integration. These encoders are pretrained via contrastive learning on combined large-scale datasets [36] and L-Mind to align neural features with semantic text embeddings.
@@ 2.1 Brain-supervised Generation
BCI, an emerging technology, builds direct communication between the brain and devices by decoding neural signals [37, 38]. Advances in machine learning have improved its accuracy, enabling brain-guided generative methods for visual content creation. Several recent methods integrate neurophysiological data (e.g., fMRI, EEG, or fNIRS) with generative models [39, 22, 40]. For instance, CMVDM aligns fMRI features with semantics for image synthesis [41], and the MindEye series further lifts the resolution of generated images from decoded fMRI [42, 43]. OneLLM leverages the large fMIR dataset for multimodal alignment in a multimodal large language model [16]. DreamDiffusion produces images from EEG via temporal masked modeling [22]. EEG2Video extends the idea to dynamic video content [44]. While Davis et al. [24] initially explore brain-guided semantic image editing using a generative adversarial network, this work is limited to facial images and EEG signals. Moreover, Adamic et al. [45] reconstructs visual images from brain activity measured by fNIRS.
@@ 2.2 Instruction-based Image Editing
Recent generative models like GPT-4o [46] and Gemini [47] have evolved from basic question answering to advanced image editing by interpreting user instructions. Modern instruction-based image editing agents integrate multimodal inputs, including text, images, and videos, to accurately identify and apply visual edits [13]. Leveraging learned multimodal representations, these agents interpret instructions from input, localize relevant regions, and perform targeted modifications [48, 49, 50]. Recent approaches, such as InstructPix2Pix [1], UltraEdit [2], MagicBrush [51], MIGE [52], and ACE [53] improve region-specific edits guided by natural language prompts. Speech-driven image editing [14] was also explored, highlighting the feasibility of hands-free interaction but still limited by linguistic expressiveness in recorded speech.
Despite these advancements, achieving efficient, delicate prompt-free image editing remains challenging. Our work addresses this gap, exploring neural-signal-driven editing agents to decode cognitive intent directly for image manipulation, significantly enhancing accessibility and interaction efficiency.
@@ 3.1 Data Collection
We collect 23,928 editing samples (22,728 training, 1,200 testing) from 12 participants using the setup depicted in Fig. 2. Participants wear our multimodal sensor while viewing image-text pairs sourced from SEED-Data-Edit [54] on a 25-inch monitor (resolution: $1980\times 1080$). The measured EEG, fNIRS, and PPG physiological signals are streamed in real-time via Bluetooth 5.3, synchronized and aligned via lab streaming layer by the proprietary Lab Recorder software [55]. Participants simultaneously read displayed editing instructions aloud, providing audio speech signals.
@@ 3.2 Data Preprocessing
PPG and motion Four-channel PPG signals (735 nm, 850 nm) are averaged per hemisphere via adaptive average pooling and filtered (0.5–4 Hz) to extract cardiac-related hemodynamic signals that reflect heart rate variability. Motion data from a six-axis sensor (12.5 Hz) capturing triaxial linear acceleration and angular velocity characterizes head movements. See supplement for more details.
@@ 4.1 Cross-Scale State Space Encoding
CS3 encoder extracts multi-scale features from diverse signals using an adaptive feature pyramid. To further capture dynamic spatio-temporal patterns beyond fixed pyramid, CS3 uses a structured state space model (S3M)  [56] for efficient long-sequence encoding with linear complexity. To manage cost, it uses a cross-feature mechanism that separately encodes temporal and channel information.
@@ 4.2 Dynamic Gated Multimodal Fusion
We calculate instance-wise and layer-wise mean $\mu$ and variance $\sigma$ from input content embedding $\mathbf{X}\!\in\!\mathbb{R}^{C^{\prime}\times L^{\prime}_{X}}$ and condition embedding $\mathbf{Y}\!\in\!\mathbb{R}^{C^{\prime}\times L^{\prime}_{Y}}$ for further fusion into $\tilde{\mathbf{H}}\!\in\!\mathbb{R}^{C^{\prime}\times L^{\prime}_{X}}$ while emphasising informative channels and suppressing noise:
@@ 4.3 Conditional Diffusion
The fused latent representation conditions a DiT backbone [57] for image editing. The DiT model accepts the encoded input image $\mathbf{I}$ and fused latent feature $\tilde{\mathbf{h}}$ and outputs the edited image aligned with the semantic intention via fine-tuning.
@@ 4.4 Pre-training and Finetuning
Pretraining. Signal encoders are pretrained to align with semantic embeddings using large-scale cognitive datasets [14] and L-Mind. CS3 encoders (EEG + PPG and fNIRS + Motion, respectively) are aligned to frozen text embeddings via symmetric NT-Xent loss:
@@ 5 Experiment
To answer the research questions (RQs) asked in Sec. 1, we conduct a comprehensive evaluation of LoongX’s neural-driven image editing capabilities on the L-Mind test set. The experimental setup, error metrics [51], and implementation details are described in Sec. A.4.1. We choose OminiControl [49] as our baseline, which supports the text-conditioned image-editing based on DiTs.
@@
5.1 Reliability of Neural Signals
As shown in Table 1, neural-signal-only LoongX outperforms the text-based OminiControl baseline in semantic discriminability (CLIP-I: $0.6605$ vs. $0.6558$) and robustness (DINO: $0.4812$ vs. $0.4636$), which demonstrates the potential of neural signals as a standalone modality that carries semantic information for image editing. The slightly higher L1 and L2 errors indicate better preservation of semantic fidelity over pixel-level accuracy. Combining speech cues with neural signals boosts semantic alignment, reaching a peak CLIP-T score of 0.2588 and demonstrating their joint effectiveness in capturing nuanced user intentions for hands-free image editing.
@@ 5.2 Ablation Studies on Modality Contribution
Modality contributions are compared in Fig. 5. EEG signals alone enable basic high-level semantic editing, supported by the semantic discriminability of the extracted features (CLIP-I: $0.5457$). Integrating fNIRS significantly improves feature robustness (DINO: from $0.2963$ to $0.4811$), highlighting the complementary nature of hemodynamic responses in enhancing signal completeness and structural fidelity. Including PPG and Motion improves global physiological awareness and indicates sensitivity to subtle engagement patterns (e.g., heart rate and user movements) that express editing intent. They both contribute to features’ robustness and completeness to ensure stable CLIP-T score gains.
We show contribution of each individual EEG channel in Fig. 5, where each channel corresponds to a specific scalp region as detailed in Table 3. The occipital cortex channel (Oz), which is in charge of visual processing, emerges as dominant in global editing effect (CLIP-I: $0.6619$) and robustness (DINO: $0.4873$) to finer details, affirming its critical role in basic visual perception and processing tasks. Conversely, the frontopolar cortex (Fpz) provides superior semantic alignment (CLIP-T: $0.2481$), consistent with its association with more complex cognitive processes. Specifically, Fpz provides decision control and attention regulation compared with basic visual perception provided by Oz, which precisely confirms the discovery patterns in medical anatomy. This channel-specific analysis provides insights valuable for targeted applications or constrained hardware settings.
Figure 4: Evaluation of different signals combined by our DGF module.
@@
The analysis of text and neural-driven image editing are shown in Fig. 6, neural signals (N) demonstrate superior performance in global texture editing, with higher CLIP-I scores highlighting their strong visual and structural consistency. Neural signals also outperform other modalities in object editings, (e.g.,object removal), reflecting their strength in conveying intuitive intent, though they remain limited in handling complex semantics. Text instructions (T) are inherently stronger in high-level semantic tasks (e.g., image restoration), which indiciates their advantage in semantic alignment. Combined neural and speech (N+S) signals achieve the highest semantic alignment (CLIP-T: $0.2588$), showcasing the superior effectiveness of hybrid conditioning in capturing complex user intentions.

### v3

@@ Abstract
Traditional image editing typically relies on manual prompting, making it labor-intensive and inaccessible to individuals with limited motor control or language abilities. Leveraging recent advances in brain-computer interfaces (BCIs) and generative models, we propose LoongX, a hands-free image editing approach driven by multimodal neurophysiological signals. LoongX utilizes state-of-the-art diffusion models trained on a comprehensive dataset of 23,928 image editing pairs, each paired with synchronized electroencephalography (EEG), functional near-infrared spectroscopy (fNIRS), photoplethysmography (PPG), and head motion signals that capture user intent. To effectively address the heterogeneity of these signals, LoongX integrates two key modules. The cross-scale state space (CS3) module encodes informative modality-specific features. The dynamic gated fusion (DGF) module further aggregates these features into a unified latent space, which is then aligned with edit semantics via fine-tuning on a diffusion transformer (DiT). Additionally, we pre-train the encoders using contrastive learning to align cognitive states with semantic intentions from embedded natural language. Extensive experiments demonstrate that LoongX achieves performance comparable to text-driven methods (CLIP-I: 0.6605 vs. 0.6558; DINO: 0.4812 vs. 0.4636) and outperforms them when neural signals are combined with speech (CLIP-T: 0.2588 vs. 0.2549). These results highlight the promise of neural-driven generative models in enabling accessible, intuitive image editing and open new directions for cognitive-driven creative technologies. The code and dataset are released on the project website: https://loongx1.github.io.
@@ 1 Introduction
Image editing involves manipulating digital visuals to achieve desired effects, significantly impacting fields like advertising, entertainment, and scientific visualization (Brooks et al., 2023). Traditionally, this task demands extensive manual effort and technical expertise. Advances in generative models have streamlined instruction-based image editing through automated pipelines (Zhao et al., 2024a; Wang et al., 2024a; OpenAI, 2025). Nevertheless, these methods still heavily depend on intensive user inputs, such as text prompts (Hertz et al., 2023; Sheynin et al., 2024), visual references like masks or sketches (Chang et al., 2023; Zeng et al., 2022), and physical operations like dragging (Shi et al., 2024; Shin et al., 2024; Liu et al., 2024d; Jiang et al., 2025). Such reliance limits efficiency and accessibility, especially for users with motor or communication impairments.
To address these challenges, alternative input modalities have been explored (Shuai et al., 2024; Jiang et al., 2021; Yang et al., 2024) for image editing. Among these, brain–computer interfaces (BCIs) provide a promising possibility with their recent advancement in hardware precision (Han et al., 2024; Allen et al., 2022). Starting from early attempts in passive tasks such as mental state recognition (Wang et al., 2024b) and neural activity analysis (Duan et al., 2023; Benchetrit et al., 2024), BCIs have begun to be involved in more active generative tasks such as neural-driven chat (Baradari et al., 2025) and visual content creation (Bai et al., 2024; Liu et al., 2024c).
However, existing approaches remain limited to the use of single-modality data such as electroencephalography (EEG) (Davis et al., 2022; Bai et al., 2024) or functional magnetic resonance imaging (fMRI) (Huo et al., 2024), which is insufficient to capture nuanced user intentions for enabling complex editing tasks. In practice, physiological signals from different modalities can offer complementary insights into cognitive states such as attention, motivation, and emotional regulation (Cai et al., 2025; Brake et al., 2024; Li et al., 2024c; Yu et al., 2014; Si et al., 2023; Arsalan and Majid, 2021; Tazarv et al., 2021), underscoring the need for multimodal neural information integration.
@@
To answer these questions, we construct L-Mind, a comprehensive multimodal dataset comprising 23,928 image pairs with synchronously collected EEG, functional near infrared spectroscopy (fNIRS) (Cao et al., 2021), photoplethysmography (PPG) (Han et al., 2020), head motion, and speech signals from 12 participants conceiving image editing tasks. Captured using a wireless, lightweight BCI system that supports unconstrained head movements and speech (Li et al., 2024a), L-Mind offers higher ecological validity under natural real-world conditions and supports robust training of brain-supervised generative models.
Building on L-Mind, we propose LoongX, a hands-free image editing approach that innovatively integrates the proposed multimodal neural signal fusion strategy with a diffusion transformer (DiT) to translate neural intent into image edits. Unlike prior single-modal methods, LoongX integrates EEG, fNIRS, PPG, and head motion signals, extracting explicit user intentions from EEG signals across multiple scalp regions, incorporating cognitive load and emotional valence data from fNIRS, and capturing stress and engagement indicators through PPG and motion signals. We introduce two new modules to manage diverse multimodal input: a cross-scale state space (CS3) encoder for robust feature extraction and a dynamic gated fusion (DGF) module for comprehensive multimodal integration. These encoders are pretrained via contrastive learning on combined large-scale datasets and L-Mind to align neural features with semantic text embeddings.
@@ 2.1 Brain-supervised Generation
As an emerging technology, brain–computer interface (BCI) builds direct communication between the brain and devices by decoding neural signals (Edelman et al., 2025; Eldawlatly, 2024). Advances in machine learning have improved its accuracy, enabling brain-guided generative methods for visual content creation. Several recent methods integrate neurophysiological data (e.g., fMRI, EEG, or fNIRS) with generative models (Takagi and Nishimoto, 2023a; Bai et al., 2024; Tripathy et al., 2021). For instance, CMVDM aligns fMRI features with semantics for image synthesis (Zeng et al., 2024), and the MindEye series further lifts the resolution of generated images from decoded fMRI (Scotti et al., 2023, 2024). OneLLM leverages the large fMIR dataset for multimodal alignment in a multimodal large language model (Han et al., 2024). DreamDiffusion produces images from EEG via temporal masked modeling (Bai et al., 2024). EEG2Video extends the idea to dynamic video content (Liu et al., 2024b). While Davis et al. (Davis et al., 2022) initially explore brain-guided semantic image editing using a generative adversarial network, this work is limited to facial images and EEG signals. Moreover, Adamic et al. (Adamic et al., 2024) reconstructs visual images from brain activity measured by fNIRS.
@@ 2.2 Instruction-based Image Editing
Recent generative models like GPT-4o (OpenAI, 2024) and Gemini (Team et al., 2023) have evolved from basic question answering to advanced image editing by interpreting user instructions. Modern instruction-based image editing agents integrate multimodal inputs, including text, images, and videos, to accurately identify and apply visual edits Shuai et al. (2024). Leveraging learned multimodal representations, these agents interpret instructions from input, localize relevant regions, and perform targeted modifications Zhang et al. (2025b); Tan et al. (2024, 2025). Recent approaches, such as InstructPix2Pix (Brooks et al., 2023), UltraEdit (Zhao et al., 2024a), MagicBrush (Zhang et al., 2023), MIGE (Tian et al., 2025), and ACE (Mao et al., 2025a) improve region-specific edits guided by natural language prompts. Speech-driven image editing (Jiang et al., 2021) was also explored, highlighting the feasibility of hands-free interaction but still limited by linguistic expressiveness in recorded speech.
Despite these advancements, achieving efficient, delicate, prompt-free image editing remains challenging. Our work addresses this gap, exploring neural-signal-driven editing agents to decode cognitive intent directly for image manipulation. It is believed that the findings in this work can significantly enhance accessibility and interaction efficiency in future BCI-enabled applications, particularly benefiting individuals with physical disabilities.
@@ 3.1 Data Collection
We collect 23,928 editing samples (22,728 training, 1,200 testing) from 12 participants using the setup depicted in Fig. 2. Participants wear our multimodal sensor while viewing image-text pairs sourced from SEED-Data-Edit Ge et al. (2024) on a 25-inch monitor (resolution: $1980\times 1080$). The measured EEG, fNIRS, and PPG physiological signals are streamed in real-time via Bluetooth 5.3, synchronized and aligned via lab streaming layer by the proprietary Lab Recorder software (Stenner et al., 2023). Participants simultaneously read displayed editing instructions aloud, providing audio speech signals.
@@ 3.2 Data Preprocessing
PPG and motion Four-channel PPG signals (735 nm, 850 nm) are averaged per hemisphere via adaptive average pooling and filtered (0.5–4 Hz) to extract cardiac-related hemodynamic signals that reflect heart rate variability. Motion data from a six-axis sensor (12.5 Hz), capturing triaxial linear acceleration and angular velocity, characterizes head movements. See supplement for more details.
@@ 4.1 Cross-Scale State Space Encoding
CS3 encoder extracts multi-scale features from diverse signals using an adaptive feature pyramid. To further capture dynamic spatio-temporal patterns beyond the fixed pyramid, CS3 uses a structured state space model (S3M)  Gu et al. (2022) for efficient long-sequence encoding with linear complexity. To manage cost, it uses a cross-feature mechanism that separately encodes temporal and channel information.
@@ 4.2 Dynamic Gated Multimodal Fusion
We calculate instance-wise and layer-wise mean $\mu$ and variance $\sigma$ from input content embedding (e.g., EEG) $\mathbf{X}\!\in\!\mathbb{R}^{C^{\prime}\times L^{\prime}_{X}}$ and condition embedding (e.g., PPG) $\mathbf{Y}\!\in\!\mathbb{R}^{C^{\prime}\times L^{\prime}_{Y}}$ for further fusion into $\tilde{\mathbf{H}}\!\in\!\mathbb{R}^{C^{\prime}\times L^{\prime}_{X}}$ while emphasising informative channels and suppressing noise:
@@ 4.3 Conditional Diffusion
The fused latent representation conditions a DiT backbone Peebles and Xie (2023) for image editing. The DiT model accepts the encoded input image $\mathbf{I}$ and fused latent feature $\tilde{\mathbf{h}}$ and outputs the edited image aligned with the semantic intention via fine-tuning.
@@ 4.4 Pre-training and Finetuning
Pretraining. Signal encoders are pretrained to align with semantic embeddings using large-scale cognitive datasets (Nieto et al., 2022; Ning et al., 2024) and L-Mind. CS3 encoders (EEG + PPG and fNIRS + Motion, respectively) are aligned to frozen text embeddings via symmetric NT-Xent loss:
@@ 5 Experiment
To answer the research questions (RQs) asked in Sec. 1, we conduct a comprehensive evaluation to validate the effectiveness of LoongX on the test set of L-Mind. This section first describes the experimental setup, evaluation metrics, and implementation details, and presents results of comprehensive quantitative evaluations, detailed breakdown analyses, and qualitative assessments.
5.1 Experimental Setup
Implementation Details.
All models are trained on eight NVIDIA H100 GPUs. Text prompts are embedded by T5-XXL Raffel et al. (2020) and CLIP Radford et al. (2021); neural signal streams are encoded by the proposed CS3. Unless stated otherwise, EEG montage (Fz, Fp2, O2, Pz, Cz) is sampled at 256 Hz and down-sampled to 32 Hz after band-pass filtering. Inference runs at 8 steps with classifier-free guidance $w=4$. We choose OminiControl (Tan et al., 2024) as our baseline as it supports the text-conditioned image-editing based on DiTs. We also implement LoongX using only neural signals (EEG, fNIRS, PPG and Motion) and using both text prompts and neural signals. We load the pretrained weights from FLUX.1-dev [^1] https://huggingface.co/black-forest-labs/FLUX.1-dev and use low-rank approximation (LoRA) for fine-tuning (learning rate $1.0$, weight decay $0.01$).
Evaluation Metrics.
We mainly use five metrics for quantitative assessment following (Zhang et al., 2023):
  1. 1)
    L1 Distance (Mean Absolute Error): Calculating the average absolute difference between corresponding pixels in edited and ground truth tar images.
  2. 2)
    L2 Distance (Mean Squared Error): Computing the average squared difference between pixels. Penalizes large errors more heavily than L1.
  3. 3)
    CLIP-I Score: Evaluating semantic similarity between model-edited images and ground truth target images, which focuses on global semantics of editing results.
  4. 4)
    DINO Score: Assessing feature similarity between editing results and ground truth. Compared with CLIP, it is believed that DINO features capture fine-grained structural similarity that correlates with perceived preservation of identity, pose, and local geometry Zhou et al. (2024).
  5. 5)
5.2 Reliability of Neural Signals
@@
As shown in Table 1, neural-signal-only LoongX outperforms the text-based OminiControl baseline in semantic discriminability (CLIP-I: $0.6605$ vs. $0.6558$) and robustness (DINO: $0.4812$ vs. $0.4636$), which demonstrates the potential of neural signals as a standalone modality that carries rich semantic information for image editing. The slightly higher L1 and L2 errors indicate better preservation of semantic fidelity over pixel-level accuracy. Combining speech cues with neural signals boosts semantic alignment, reaching the highest CLIP-T score of 0.2588 and demonstrating their joint effectiveness in capturing nuanced user intentions for hands-free image editing.
Figure 4: Evaluation of different signal combinations on the proposed DGF module.
@@ 5.3 Ablation Studies on Modality Contribution
Modality contributions are compared in Fig. 5. EEG signals alone enable basic high-level semantic editing, supported by the semantic discriminability of the extracted features (CLIP-I: $0.5457$). Integrating fNIRS significantly improves feature robustness (DINO: from $0.2963$ to $0.4811$), highlighting the complementary nature of hemodynamic responses in enhancing signal completeness and structural fidelity. Including PPG and Motion improves global physiological awareness and indicates sensitivity to subtle engagement patterns (e.g., heart rate and user movements) that express editing intent. They both contribute to the features’ robustness and completeness to ensure stable CLIP-T score gains.
We show the contribution of each individual EEG channel in Fig. 5, where each channel corresponds to a specific scalp region as detailed in Table 3. The occipital cortex channel (Oz), which is in charge of visual processing, emerges as dominant in global editing effect (CLIP-I: $0.6619$) and robustness (DINO: $0.4873$) to finer details, affirming its critical role in basic visual perception and processing tasks. Conversely, the frontopolar cortex (Fpz) provides superior semantic alignment (CLIP-T: $0.2481$), consistent with its association with more complex cognitive processes. Specifically, Fpz provides decision control and attention regulation compared with basic visual perception provided by Oz, which precisely confirms the discovery patterns in medical anatomy. This channel-specific analysis provides insights valuable for targeted applications or constrained hardware settings.
@@
The analysis of text and neural-driven image editing is shown in Fig. 6. Using pure neural signals (N) is particularly effective for global texture editing, with higher CLIP-I scores highlighting their strong visual and structural consistency. Neural signals also outperform other modalities in several tasks like object removal and background blur, reflecting their strength in conveying intuitive intent, though they remain limited in handling complex semantics like text editing. Text instructions (T) are inherently stronger in high-level semantic tasks (e.g., image restoration), which indicates their advantage in describing instruction details. Combined neural and speech (N+S) signals achieve the highest semantic alignment (CLIP-T: $0.2588$), showcasing the superior effectiveness of hybrid conditioning in capturing complex user intentions. Overall, neural signals are more effective for low-level visual edits, while neural and text-based approaches each provide complementary advantages.